OpenStudy (anonymous):

dy/dx of x^2sinxtanx Ii got....2xsinx+x^2sec^2xtanx+x^2sinxsec^x ...but i'm not sure if that is correct?

OpenStudy (anonymous):

\[ y = x^{\sin x \tan x} \]So \[ \ln y =\sin x \tan x \ln x \]then try differentiation.

OpenStudy (anonymous):

its x^2 is the only exponent..not sinxtanx

OpenStudy (anonymous):

sorry about the confusion.

OpenStudy (anonymous):

\[x ^{2} \]